Fermi National Laboratory
 

Operation Summary for October 22, 2010 - October 25, 2010

 

The Integrated Luminosity for the period from 10/18/10 to 10/25/10 was 51.64 pb-1.
NuMI reported receiving no protons on target during this same period due to target work.

Friday Oct 22

The day shift began with Operations monitoring the Tevatron (TeV) store 8198 (185.41E30) and the Recycler stash (154E10), with the Antiproton Source (Pbar) stacking (15.6E10), with Linac, Booster, and Main Injector (MI) operating normally, with the NuMI experiment off for target work, with the MiniBooNE and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations reported that Recycler Electron Cooling (ECool) personnel continued to investigate solenoid corrector issues. EE Support personnel investigated LCW flow interlock issues with a Meson power supply (M01D).

During the evening shift, Operations terminated store 8198 at 5:05 PM, put the TeV into HEP shot setup, and established store 8199 at 6:20 PM with an average initial luminosity (AIL) of 331E30.

Saturday Oct 23

The midnight shift began with Operations monitoring store 8199 (142.23E30) and the stash (183.7E10), with Pbar stacking (14.6E10), and with MiniBooNE taking beam.

During the midnight shift, at 5:25 AM, a Pbar Debuncher (D:Q804) power supply tripped off. Operators found the power supply without control power. Operations called in Pbar personnel to investigate. Operations established beam to T-978 at 6:15 AM. Operators added insulating gas to the TeV sector E48 kicker system.

During the day shift, Pbar personnel found and repaired a bad blower for D:Q804; stacking resumed at 8:45 AM. Operations reported that store 8199 aborted due to the A1-A2 beam valves closing. Mechanical Support personnel investigated and found a problem with the air dryer at A1. Operators prepared the Transfer Hall for access to open the separator valve. Operators started the TeV turn-on at 11 AM. Operations checked the TeV orbits, put the TeV into shot setup at 1:30 PM, and established store 8202 at 2:30 PM with an AIL of 332E30. Operations reported that a Booster vacuum burst caused a RF station (BRF18) to trip off. Booster personnel investigated because a vacuum pump kept tripping off. Operators decreased the Booster beam intensity while the vacuum recovered.

During the evening shift, Operations reported that Booster RF watchdog trips held off beam. Operators called in Booster personnel; they found a problem with the anode program and with a power supply (Noker) that was sporadically notching the second stacking pulse. Booster personnel lowered the supply voltage and everything returned to nominal.

Sunday Oct 24

The midnight shift began with Operations monitoring store 8202 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8202 at 3:15 AM. A Booster kicker (MKS09) tripped off at 3:25 AM due to a bad low-level signal chassis. Booster personnel came in and replaced the chassis. During the repair, Operations put the TeV through a dry squeeze. Operations put the TeV into shot setup at 6:15 AM and established store 8204 at 7:25 AM with an AIL of 334E30. Operations then established beam to T-978.

During the day shift, operators reported that the H- Source voltage regulation went unstable; operators recovered the source. A MI controls crate developed a power supply problem that inhibited stacking, operators replaced the supply.

During the evening shift, Operations terminated store 8204 at 8:05 PM, put the TeV into shot setup, and established store 8205 at 9:21 PM with an AIL of 330E30.

Monday Oct 25

The midnight shift began with Operations monitoring store 8205 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, Operations reported that a Switchyard power supply (S:HP3US) would not turn on for slow spill operations to T-978. Operations called in EE Support personnel to investigate. At 6:50 AM, a MI kicker (KPS5S) developed a Thyratron power supply fault; kicker personnel investigated.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will terminate store 8205 this morning and allow cryo system personnel to replace a wet engine flywheel at TeV sector B4.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/25/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory